Grade 11Mechanics

16gm Oxygen expand at constant pressure and 273.15 K to occupy double its original volume. The work done during process will be?

Devendra Verma
8 Years agoGrade 11

1 Answer

Satabdi das
7 Years ago
Since final volume is twice that of initial volume so 
W= P*(Vfinal-Vinitial) 
W=1*(2V-V)
W=V
Now no. Of moles = 16000/32= 500 
PV=nRT 
V=500*2*273=27300
Now W =27300/1000= 273kcal